OpenStudy (anonymous):

May i please get some help! I dont understand and have been struggling with this for a week and about to get kicked out for not doing work... Please help! :'(

OpenStudy (anonymous):

Simplify the following expression: (x - 8y) - (2x - 4y). If the final answer is written in the form Ax + By, what is the value of B?

OpenStudy (fibonaccichick666):

can you combine like terms?

OpenStudy (anonymous):

A little... :'(

OpenStudy (fibonaccichick666):

ok, show me what you got :)

OpenStudy (anonymous):

3x-12y??

OpenStudy (fibonaccichick666):

really good attempt, you only forgot 1 thing.

OpenStudy (anonymous):

What? I have to know...

OpenStudy (fibonaccichick666):

\[ (x - 8y) \color\red{-} (2x - 4y)= (x - 8y) [\color\red{-1*} (2x - 4y)]\]

OpenStudy (fibonaccichick666):

You forgot to distribute the minus sign

OpenStudy (fibonaccichick666):

try thinking of it as I wrote it above.

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

Okay... Idk... :'(

OpenStudy (fibonaccichick666):

ok, so, can you just distribute the minus sign for me here \[-(2x-4y)\]?

OpenStudy (fibonaccichick666):

watch the video if you are confused, then come back and tag me

OpenStudy (anonymous):

For some reason its not coming up @FibonacciChick666

OpenStudy (fibonaccichick666):

hmm... try google ing "khanacademy: simplifying parentheses" it is the second link

OpenStudy (anonymous):

Okay... Wait i brb...

OpenStudy (anonymous):

Sorry i dont get it... Could you just help me step by step? PLease!

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (anonymous):

@Ashy98

OpenStudy (anonymous):

@juana02

OpenStudy (anonymous):

@flight98

OpenStudy (anonymous):

@piercetheveil47

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@paki

OpenStudy (ashy98):

x-8y-2x+4y = -x-4y A=-1,B=-4

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

@Directrix

OpenStudy (ashy98):

did you watch that video @FibonacciChick666 put ?

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

Its okay i got it... May you guys help with another?

OpenStudy (fibonaccichick666):

No? .

OpenStudy (anonymous):

@ckellar2

OpenStudy (anonymous):

Okay?.... I mean from everyone else besides you @FibonacciChick666

OpenStudy (anonymous):

whts next?

OpenStudy (anonymous):

What ya mean??

OpenStudy (anonymous):

nvm

OpenStudy (jhannybean):

Could you show us how you solved it?

OpenStudy (anonymous):

Umm... I didnt get it... So I just skipped it... Its taking me too long...

OpenStudy (jhannybean):

That's not how you're going to learn :(

OpenStudy (anonymous):

i was goiing to go back t it later...

OpenStudy (jhannybean):

So you have ` (x - 8y) - (2x - 4y)` what is -1 \(\cdot\) (2x) = ?

OpenStudy (anonymous):

It combined i guess?

OpenStudy (jhannybean):

what is a number multiplied by -1 ......

OpenStudy (anonymous):

The same when it was multiplied??

OpenStudy (jhannybean):

No. When you multiply a number, x by -1, the number becomes negative. -x.

OpenStudy (jhannybean):

So according to this understanding, if you multiply a number, 2x, by -1, what does it become? \((-1)(2x) = ~? \)

OpenStudy (anonymous):

Oh... Okay.

OpenStudy (anonymous):

-2x??

OpenStudy (jhannybean):

Good job.

OpenStudy (anonymous):

YAY!! :)

OpenStudy (jhannybean):

Now when you multiply 2 negatives together, you get a + number. i.e. \((-1)(-1)= +1\). So what is \(-1)(-4y)=~?\)

OpenStudy (anonymous):

5?

OpenStudy (jhannybean):

How did you get 5.

OpenStudy (anonymous):

Since it turns into i positive i figured it would add instead of subtract... Am i right?

OpenStudy (jhannybean):

\[(-1)(-4y) \ne 5\] try again.

OpenStudy (jhannybean):

No, parenthesis always suggest multiplication.

OpenStudy (anonymous):

3?

OpenStudy (jhannybean):

look at the example I gave you with the 2 negative 1's. You are not actively paying attention right now and it is reflecting through your understanding. It seems more as if you are guessing rather than even bothering to read my posts.

OpenStudy (anonymous):

I am reading your post!! Sorry im a little slow...

OpenStudy (anonymous):

:'(

OpenStudy (usukidoll):

ok.. what is 4 x 1?

OpenStudy (usukidoll):

?

OpenStudy (anonymous):

4...

OpenStudy (usukidoll):

alright good... now what (-4)(-y) negative x negative = positive

OpenStudy (anonymous):

4y

OpenStudy (usukidoll):

yes ^_^

OpenStudy (anonymous):

Okay... YAY!

OpenStudy (usukidoll):

so back to this (x - 8y) - (2x - 4y) (x-8y-2x+4y) rewrite it... combing all x's and y's and then do the computation

OpenStudy (usukidoll):

combine all the x's first then y's and then simplify

OpenStudy (anonymous):

Okay... 3x-12y....

OpenStudy (usukidoll):

ummm... we are just adding and subtracting

OpenStudy (usukidoll):

(x-8y-2x+4y) ok take this and rearrange it... have all x's by itself and all y's by itself

OpenStudy (anonymous):

x-2x-8y+4y???

OpenStudy (jhannybean):

\(\checkmark\)

OpenStudy (usukidoll):

ok now solve x and y

OpenStudy (usukidoll):

as in what is x-2x what is -8y+4y

OpenStudy (anonymous):

Umm... I am guessing 1x 14y?

OpenStudy (anonymous):

@UsukiDoll ?? Ya there? Sorry i took long...

OpenStudy (usukidoll):

ya and it's wrong

OpenStudy (anonymous):

Oh... :'( I will never get it...

OpenStudy (usukidoll):

think of it as ... for x... I am selling a brownie for $2 but you only have $1, so how much do you owe me?

OpenStudy (anonymous):

$1 back...

OpenStudy (usukidoll):

yeah you owe me $1 so , you're -1 in the red.. so that's -x

OpenStudy (anonymous):

Okay...

OpenStudy (usukidoll):

again with this one -8y+4y I'm selling a plate lunch for $8, but you only have $4 ... you owe me...

OpenStudy (anonymous):

$4...

OpenStudy (usukidoll):

yes

OpenStudy (usukidoll):

-x-4y

OpenStudy (anonymous):

OKay... How about this one?

OpenStudy (usukidoll):

that's the answer

OpenStudy (anonymous):

OH!! =o

OpenStudy (anonymous):

Thank you! :)